OpenStudy (anonymous):

If sinA = 2/3 then find cosA = ? with method please...

Nnesha (nnesha):

draw a triangle  \[\large\rm sin \theta =\frac{ opposite }{ hypotenuse } ~~~ \cos \theta=\frac{ adjacent }{ hypotenuse } ~~\tan \theta =\frac{ opposite }{ adjacent }\]

Nnesha (nnesha):

 apply Pythagorean theorem to find side of right triangle \[\huge\rm a^2+b^2=c^2\]